OpenStudy (anonymous):

Simplify each exponential expression using the appropriate rule of exponents (7 (123456.789)^4)^0

OpenStudy (campbell_st):

well this is the power of a power law which means the outer most power ), operates on each part of the term in the brackets. and the rule is multiply the powers \[(ab^n)^m = a^m \times b^{n \times m}\] then you need to know about the power of zero which is anything to a power of zero is equal to 1. \[x^0 = 1\] hope it helps

OpenStudy (anonymous):

so all of this equals zero?

OpenStudy (anonymous):

1* i mean